# The Simons Observatory: Validation of reconstructed power spectra from simulated filtered maps for the Small Aperture Telescope survey

### v1

@@ 1 Introduction
A detection of primordial tensor perturbations, or a sufficiently tight upper bound on $r$, could elucidate many of the remaining questions about the beginning of our Universe, allowing us to discriminate between different models. Cosmic inflation, which posits a short phase of exponentially accelerated expansion at early times, is the most widespread of these models [e.g., 6, 7, 8]. Within this paradigm, primordial perturbations are seeded by quantum fluctuations in the spacetime metric that are expanded to macroscopic scales during the inflationary period [e.g., 9, 10, 11, 12]. The presence of both scalar and tensor fluctuations (i.e. PGWs), is a key prediction of this theory [e.g., 13]. Current and upcoming CMB experiments such as the BICEP Array [14], SPIDER [15], CLASS [16], AliCPT-1 [17], CMB-S4 [18], and LiteBIRD [19] are targeting or will target the measurement of these elusive primordial $B$-modes in the next decade.
Polarised foreground emission from our galaxy is a substantial complication for measuring $r$ since it produces a $B$-mode signal much larger than the one from primordial $B$-modes. Thermal dust and synchrotron emission are the two main polarised components that could pose a problem [e.g., 20]. Their $B$-mode emission would be equivalent to a tensor-to-scalar ratio much higher than $r\sim 0.01$ [21], the target of many upcoming CMB experiments. By exploiting the difference in spectral distributions between the CMB and foregrounds, we can clean the former, using methods collectively known as component separation [22]. However, small residuals left after cleaning may still pose a problem for measuring small values $r\sim 0.001$ in the near future, since the residuals can be as large as the primordial signal we want to measure in the first place.
@@
Ground-based observatories contend with multiple sources of low-frequency noise that can affect measurements of the primordial $B$-mode signal. For example, temporal and spatial fluctuations in the emission of the atmosphere due to turbulence and uneven distribution of water vapour, which creates non-trivial correlations between detectors [e.g., 25], as well as by intensity-to-polarisation (I-to-P) leakage due to factors such as thermal drifts, optical mismatch, etc. When doing detector pair differencing with orthogonal antennas, bandpass or beam mismatches arise [e.g., 26, 27]. To address these difficulties, the SO-SATs modulate the polarisation signal at frequencies where detector white noise dominates, above atmospheric fluctuations and thermal drifts, using a cryogenic HWP rotating around $f\approx 2$ Hz [28]. This technique also eliminates the need for detector pair differencing, allowing each detector to independently measure the three Stokes parameters $(I,Q,U)$ [29]. Previous experiments that have used HWP modulation include MAXIPOL [30], ABS [31], and Polarbear [32], while CLASS uses a different modulation technology, the variable-delay polarization modulator [VPM, 33]. A demodulation at the modulating fourth harmonics of the HWP rotating frequency, followed by a low-pass filter, is the standard procedure to recover the polarised signal [e.g., 34]. A further optional high-pass filter can be applied to the polarised timestreams to suppress the remaining low-frequency $1/f$ noise. The impact of all this filtering must be accounted for when generating maps from the filtered timestreams, or the associated bias to the maps must be taken into account and corrected when analysing them [e.g., 30, 35, 36, 37, 38]. General filtering to remove low-frequency sources of noise is a standard practice in CMB science, even without polarised modulation, and these effects therefore need to be accounted for in cosmological analyses [e.g., 39, 40, 41, 42, 43, 44].
@@ 2.1 Filter$+$bin mapmaking
Ground-based CMB telescopes typically observe the sky at fixed elevation, scanning back and forth in azimuth, and taking samples at a regular rate with thousands of detectors in the telescope focal plane. While each individual sample is noise-dominated, the sky signal can be teased out by accumulating samples collected over time from the same sky coordinates in a process called “mapmaking” [e.g., 45]. The relation between a given measured timestream and the true sky is given by
@@
This process would be significantly simpler in the presence of a diagonal noise covariance $\bm{\mathrm{N}}$, corresponding to the case of white noise and uncorrelated detectors. In this case, the LHS matrix is also diagonal across pixels and can be expressed as $N_{\rm pixels}$ “binned” matrices of shape $N_{\rm comps}\times N_{\rm comps}$ (or equivalently one $N_{\rm comps}N_{\rm pixels}\times N_{\rm comps}N_{\rm pixels}$ matrix), which can be easily inverted. The right-hand side (RHS) $\bm{\mathrm{P}}^{T}\bm{\mathrm{N}}^{-1}\bm{d}$ is also projected to map space, being a vector with shape $1\times N_{\rm comps}N_{\rm pixels}$.
Although modulating and demodulating the signal (e.g., using a continuously rotating HWP) should whiten the noise, this idealised situation is never exactly realised. We may nevertheless follow this approach and correct for the impact of the associated approximations a posteriori. In this case, it is common to work with filtered timestreams, in order to deal with systematics and inverse-variance-weight the data. In other words, we apply a set of filters $\bm{\mathrm{F}}$ to the data, and project, or “bin”, the filtered data $\bm{d}^{\prime}\equiv\bm{\mathrm{F}}\bm{d}$ before solving the mapmaking equation under the assumption of a diagonal noise covariance. The maps resulting from this procedure (commonly called “filter$+$bin” mapmaking), and their directly estimated power spectra, will be biased by an unknown multiplicative transfer function (TF). This TF can then be estimated and corrected via simulations, as described in Sections 2.2.1 and 2.2.2. Our goal is to use a relatively fast method that works on the principle of the TF and some approximations.
@@ 2.2 From maps to spectra
$B^{\alpha\beta b}_{\alpha^{\prime}\beta^{\prime}\ell^{\prime}}\equiv\sum_{\alpha^{\prime\prime}\beta^{\prime\prime}b^{\prime\prime}}({\cal W}^{-1})_{\alpha^{\prime\prime}\beta^{\prime\prime}b^{\prime\prime}}^{\alpha\beta b}\sum_{\ell^{\prime\prime}\in b^{\prime\prime}}v_{b^{\prime\prime}}^{\ell^{\prime\prime}}\,X_{\alpha^{\prime}\beta^{\prime}\ell^{\prime}}^{\alpha^{\prime\prime}\beta^{\prime\prime}\ell^{\prime\prime}},$ (2.11)
as opposed to equation (2.9). Since we cannot calculate $X$ analytically, we must resort to a simulation-based approach in order to estimate the BPWFs. The next two sections present two different approaches to achieve this.
@@ 2.2.1 Mode coupling $-\ell$
Figure 1: Comparison of the reconstructed band power window functions using the $-\ell$ simulation-based method described in Section 2.2.1. The dashed black line shows the analytical NaMaster-calculated BPWF for $EE\rightarrow EE$ and a partial sky mask at $N_{\rm side}=64$. The dotted red line shows the simulation-based BPWFs averaged across 100 simulations per $\ell$. The red shaded area represents the standard deviation across the 100 simulations divided by $\sqrt{100}$. For this example, we use the mask constructed from the hits map in Figure 2. The other BPWFs look qualitatively similar. We remind the reader this is an example of a test case with no time-domain filtering.
We can partially validate this method in a simple way, by considering the limit in which no filtering operations are applied to the data beyond masking. In this case, the MCM and the BPWFs can be estimated analytically using the standard pseudo-$C_{\ell}$ algorithm via NaMaster [56]. We can then compare this analytical solution with the estimate from $\delta$-like simulations. Figure 1 shows the BPWFs estimated through these two methods, with the analytical solution as the dashed black line, and the simulation-based estimate as the dotted red line, using 100 simulations per $\ell$. Although the simulation-based estimate is intrinsically noisy, the differences between them are consistent within the sample variance due to the limited number of simulations.
The main drawback of this method is the large number of simulations needed to reconstruct the BPWF. As we will show in Section 4.1, an accurate estimate of the BPWFs, particularly accounting for $E/B$ leakage, may require $\mathcal{O}(10^{5})$ simulations, on which expensive filtering operations must be carried out, in order to reconstruct the first ${\sim}200$ multipoles for SO-SAT. This is almost impossible to achieve when doing end-to-end simulated filtering. This process may be accelerated through the use of observation matrices (e.g., as done by the BICEP collaboration [42]), as long as constructing this matrix is feasible given available computing resources. This may pose challenges for experiments covering large sky areas, such as SO-SAT, or reaching few-arcminute resolutions. This motivates exploring alternative approximate but less computationally demanding methods.
@@ 2.2.2 Using a transfer function
    We average both sets of power spectra over a number of simulations ($100$ in our case). Note that we may calculate the unfiltered power spectra $\tilde{C}^{{\rm PL},(\alpha\beta)(\gamma\delta)}_{b}$ analytically because the impact of masking can be calculated exactly using NaMaster.
@@
With the TF estimated as we just described, we can construct our model for the mode-coupling matrix of the band powers by combining it with the mask-only mode-coupling matrix (equation (2.12)). The method then proceeds as it does for the standard pseudo-$C_{\ell}$ approach: we bin and invert this mode-coupling matrix, and use it to estimate the mode-decoupled band powers and the associated BPWFs. Our implementation of this method [^5] https://github.com/simonsobs/SOOPERCOOL v0.0 includes all the cross-spectra of the $IEB$ fields. As stated, this paper will only focus on polarisation and ignore intensity.
@@ 3.1 Input simulated maps
All of our simulated maps are in healpix pixels [57], produced at single float precision. For a comparison between the numerical $-\ell$ and TF methods, shown in Section 4.1, we use maps at resolution $N_{\rm side}=128$. The limited resolution is necessary because of the observation matrix (this is explained in Section 3.3.2). For the TF method results at full resolution and multipole range, shown in Section 4.2, we use maps at resolution $N_{\rm side}=512$. All the angular power spectra in this work are calculated with NaMaster, which corrects for the mode coupling arising from partial sky observation. We do not use $E$/$B$-mode purification. [^6] This analysis aims at showing our results are unbiased and not at optimizing the power spectra estimators. While we could have used purification, we decided not to since the interplay between filtering and purification is the subject of ongoing work. The masks are constructed from normalised hits maps and with a C1 apodisation [58] with a $5^{\circ}$ length. [^7] See https://namaster.readthedocs.io/en/latest/api/pymaster.utils.html#pymaster.utils.mask_apodization for details.
@@ 3.2 Simulated dataset
In total, our schedule has 167 observations for a total integration time of 139 hours. Each observation is almost an hour long and includes the 7 detector arrays, known as wafers, covering the focal plane of one SAT, for a total of $\sim 6000$ detectors per each f090/f150 band (however we use only one-eighth of these as described below). The normalised hits map of our full schedule is shown in Figure 2, scanning the South Galactic patch. Our schedule is run in 8 sub-schedules that produce one filter+bin map each. These maps are linearly co-added. [^10] Since the filtering and mapmaking operations are linear, coadding individual sub-schedule filter+bin maps yields the same map as applying the filter+bin mapmaker to the total schedule. In particular, 5 sub-schedules correspond to observations when the South patch is rising through the sky (equivalent to 102 observations and 93.5 hours) and 3 sub-schedules when the South patch is setting in the sky (equivalent to 65 observations and 45.5 hours). This separation into sub-schedules is useful when defining splits for simulating a null test or trying to estimate cross-spectra without noise bias.
@@ 3.3 Using toast
  • A HWP synchronous signal filter is used to subtract the HWP synchronous signal introduced in the modulation of the TOD. This is a sines/cosines template fitted to data as a function of HWP angle and subtracted.
@@
  • A ground filter is used to subtract the horizontal-coordinate-fixed contamination from the ground. This is a Legendre polynomial of order 10 fitted to data as a function of azimuth and subtracted.
  • A high-pass filter is used to suppress the impact of low-frequency contamination such as the atmosphere and detector thermal drifts, among others. This is a Legendre polynomial of order 3 fitted to the data as a function of time and subtracted. This and the ground filter are applied simultaneously since they are templates deprojected in one matrix operation.
In Appendix A, we illustrate the effect of individual filters. Finally, the TOD are binned into healpix $N_{\rm side}=512$ maps according to their pointing. We make some approximations to ease the computing cost of running these simulations for this work. We use a 40 Hz sampling rate for scanning (real observations scan at 200 Hz) and we also “thin” the focal plane by a factor of 8, i.e., only one out of every eight detectors is simulated. In Appendix B we discuss this in detail, showing how the impact is minimal while reducing the computation time by a factor of 8. The combined effect of the three data volume reductions (one day every month; 40 Hz; 1/8th of detectors) is to reduce the computational requirements of the simulations by a factor of 1200 per frequency band per SAT.

### v2

@@ 1 Introduction
A detection of primordial tensor perturbations, or a sufficiently tight upper bound on $r$, could elucidate many of the remaining questions about the beginning of our Universe, allowing us to discriminate between different models. Cosmic inflation, which posits a short phase of exponentially accelerated expansion at early times, is the most widespread of these models [e.g., 6, 7, 8]. Within this paradigm, primordial perturbations are seeded by quantum fluctuations in the spacetime metric that are expanded to macroscopic scales during the inflationary period [e.g., 9, 10, 11, 12]. The presence of both scalar and tensor fluctuations (i.e. PGWs), is a key prediction of this theory [e.g., 13]. Current and upcoming CMB experiments such as BICEP/KECK [5, 14], BICEP Array [15], SPIDER [16], CLASS [17], Polarbear [18], QUBIC [19], GroundBIRD [20], AliCPT-1 [21], CMB-S4 [22], and LiteBIRD [23] are targeting or will target the measurement of these elusive primordial $B$-modes in the next decade.
Polarised foreground emission from our galaxy is a substantial complication for measuring $r$ since it produces a $B$-mode signal much larger than the one from primordial $B$-modes. Thermal dust and synchrotron emission are the two main polarised components that could pose a problem [e.g., 24]. Their $B$-mode emission would be equivalent to a tensor-to-scalar ratio much higher than $r\sim 0.01$ [25], the target of many upcoming CMB experiments. By exploiting the difference in spectral distributions between the CMB and foregrounds, we can clean the former, using methods collectively known as component separation [26]. However, small residuals left after cleaning may still pose a problem for measuring small values $r\sim 0.001$ in the near future, since the residuals can be as large as the primordial signal we want to measure in the first place. Another relevant foreground contamination are the coherent distortions induced by the deflection of CMB photons passing through the gravitational potential wells of the cosmic large-scale structure, known as CMB weak lensing [27]. $B$-modes are sourced from primordial CMB $E$-modes at intermediate and small scales [28], effectively creating a lensing $B$-mode spectrum that will dominate over the primordial $B$-modes in all but the largest scales, depending on the value of $r$. For low enough values of $r$, this lensing-induced noise will dominate the sample variance and limit the precision of the cosmological parameters we can measure. We can mitigate this with the technique of delensing, in which we subtract a template of the realization of lensing $B$-modes, reconstructed from an estimator of the lensing potential and observations of the $E$-modes at intermediate and high resolution [29, 30].
@@
Ground-based observatories contend with multiple sources of low-frequency noise that can affect measurements of the primordial $B$-mode signal. For example, temporal and spatial fluctuations in the emission of the atmosphere due to turbulence and uneven distribution of water vapour, which creates non-trivial correlations between detectors [e.g., 33], as well as by intensity-to-polarisation (I-to-P) leakage due to factors such as thermal drifts, optical mismatch, etc. When doing detector pair differencing with orthogonal antennas, bandpass or beam mismatches arise [e.g., 34, 35]. To address these difficulties, the SO-SATs modulate the polarisation signal at frequencies where detector white noise dominates, above atmospheric fluctuations and thermal drifts, using a cryogenic HWP rotating around $f\approx 2$ Hz [36]. This technique also eliminates the need for detector pair differencing, allowing each detector to independently measure the three Stokes parameters $(I,Q,U)$ [37]. Previous experiments that have used HWP modulation include MAXIPOL [38], ABS [39], and Polarbear [40], while CLASS uses a different modulation technology, the variable-delay polarization modulator [VPM, 41]. A demodulation at the modulating fourth harmonics of the HWP rotating frequency, followed by a low-pass filter, is the standard procedure to recover the polarised signal [e.g., 42]. A further optional high-pass filter can be applied to the polarised timestreams to suppress the remaining low-frequency $1/f$ noise. The impact of all this filtering must be accounted for when generating maps from the filtered timestreams, or the associated bias to the maps must be taken into account and corrected when analysing them [e.g., 38, 43, 44, 45, 46]. General filtering to remove low-frequency sources of noise is a standard practice in CMB science, even without polarised modulation, and these effects therefore need to be accounted for in cosmological analyses [e.g., 47, 48, 49, 50, 51, 52].
@@ 2.1 Filter$+$bin mapmaking
Ground-based CMB telescopes typically observe the sky at fixed elevation, scanning back and forth in azimuth, and taking samples at a regular rate with thousands of detectors in the telescope focal plane. While each individual sample is noise-dominated, the sky signal can be teased out by accumulating samples collected over time from the same sky coordinates in a process called “mapmaking” [e.g., 53]. The relation between a given measured timestream and the true sky is given by
@@
This process would be significantly simpler in the presence of a diagonal noise covariance $\bm{\mathrm{N}}$, corresponding to the case of white noise and uncorrelated detectors. In this case, the LHS matrix is also diagonal across pixels and can be expressed as $N_{\rm pixels}$ “binned” matrices of shape $N_{\rm comps}\times N_{\rm comps}$ (or equivalently one $N_{\rm comps}N_{\rm pixels}\times N_{\rm comps}N_{\rm pixels}$ matrix with zero correlation between pixels), which can be easily inverted. The right-hand side (RHS) $\bm{\mathrm{P}}^{T}\bm{\mathrm{N}}^{-1}\bm{d}$ is also projected to map space, being a vector with shape $1\times N_{\rm comps}N_{\rm pixels}$.
Although modulating and demodulating the signal (e.g., using a continuously rotating HWP) should whiten the noise, this idealised situation is never exactly realised. We may nevertheless follow this approach and correct for the impact of the associated approximations a posteriori. In this case, it is common to work with filtered timestreams, in order to deal with systematics and inverse-variance-weight the data. In other words, we apply a set of filters $\bm{\mathrm{F}}$ to the data, and project, “bin”, or “coadd” the filtered data $\bm{d}^{\prime}\equiv\bm{\mathrm{F}}\bm{d}$ before solving the mapmaking equation under the assumption of a diagonal noise covariance. The maps resulting from this procedure (commonly called “filter$+$bin” mapmaking), and their directly estimated power spectra, will be biased by an unknown multiplicative transfer function (TF). This TF can then be estimated and corrected via simulations, as described in Sections 2.2.1 and 2.2.2. Our goal is to use a relatively fast method that works on the principle of the TF and some approximations.
@@ 2.2 From maps to spectra
$B^{\alpha\beta b}_{\alpha^{\prime}\beta^{\prime}\ell^{\prime}}\equiv\sum_{\alpha^{\prime\prime}\beta^{\prime\prime}b^{\prime\prime}}({\cal X}^{-1})_{\alpha^{\prime\prime}\beta^{\prime\prime}b^{\prime\prime}}^{\alpha\beta b}\sum_{\ell^{\prime\prime}\in b^{\prime\prime}}v_{b^{\prime\prime}}^{\ell^{\prime\prime}}\,X_{\alpha^{\prime}\beta^{\prime}\ell^{\prime}}^{\alpha^{\prime\prime}\beta^{\prime\prime}\ell^{\prime\prime}},$ (2.11)
as opposed to Equation (2.9). Here ${\cal X}$ is the binned version of $X$, in analogy to ${\cal W}$ in Equation (2.7). Since we cannot calculate $X$ analytically, we must resort to a simulation-based approach in order to estimate the BPWFs. The next two sections present two different approaches to achieve this.
@@ 2.2.1 Mode coupling $-\ell$
Figure 1: Comparison of the reconstructed band power window functions using the $-\ell$ simulation-based method described in Section 2.2.1. The dashed black line shows the analytical NaMaster-calculated BPWF for $BB\rightarrow BB$ (left hand side panel) and $EE\rightarrow BB$ (right-hand side panel) for a partial sky mask at $N_{\rm side}=64$. The dotted red line shows the simulation-based BPWFs averaged across 100 simulations per $\ell$. The red shaded area represents the standard deviation across the 100 simulations divided by $\sqrt{100}$. For this example, we use the mask constructed from the hits map in Figure 2. The other BPWFs look qualitatively similar. We remind the reader this is an example of a test case with no time-domain filtering.
We can partially validate this method in a simple way, by considering the limit in which no filtering operations are applied to the data beyond masking. In this case, the MCM and the BPWFs can be estimated analytically using the standard pseudo-$C_{\ell}$ algorithm via NaMaster [64]. We can then compare this analytical solution with the estimate from $\delta$-like simulations. Figure 1 shows the $BB\rightarrow BB$ and $EE\rightarrow BB$ BPWFs estimated through these two methods, with the analytical solution as the dashed black line, and the simulation-based estimate as the dotted red line, using 100 simulations per $\ell$. Although the simulation-based estimate is intrinsically noisy, the differences between them are consistent within the sample variance due to the limited number of simulations.
The main drawback of this method is the large number of simulations needed to reconstruct the BPWF. As we will show in Section 4.1, an accurate estimate of the BPWFs, particularly accounting for $E/B$ leakage, may require $\mathcal{O}(10^{5})$ simulations, on which expensive filtering operations must be carried out, in order to reconstruct the first ${\sim}200$ multipoles for SO-SAT. This is almost impossible to achieve when doing end-to-end simulated filtering [^5] As an illustration, it takes approximately 20 minutes to filter and bin one map as described in Section 3.3 with one node of the perlmutter computer at NERSC, using 128 cores. To reconstruct the first 200 multipoles, we would need to repeat this for $191\times 4\times 1000=764,000$ maps, assuming we do 1000 realizations $-\ell$. Often, we will need even more realizations to constrain the leakage terms. In this example, this would take $10,611$ node-days.. This process may be accelerated through the use of observation matrices (e.g., as done by the BICEP collaboration [50]), as long as constructing this matrix is feasible given available computing resources. This may pose challenges for experiments covering large sky areas, such as SO-SAT, or reaching few-arcminute resolutions. This motivates exploring alternative approximate but less computationally demanding methods.
@@ 2.2.2 Using a transfer function
The main assumption here (that coupling between different harmonic modes is driven by the sky mask, and not by filtering) is not entirely without basis. Consider the toy scenario of a one-dimensional sky mapped by a continuously scanning telescope. Filtering the resulting timestream is, in the most common cases (e.g. Fourier-space filters, polynomial filters), a translationally-invariant linear transformation. As such, it affects each Fourier mode of the timestream (and of the 1-dimensional sky) independently, without mixing them. This is similar to beam smoothing, although most filters instead affect the largest scales. This analogy breaks down for the true two-dimensional sky, especially in regions of poor cross-linking, and when filtering occurs over finite time periods (e.g. telescope subscans), breaking translational invariance. Nevertheless, it illustrates the reasons for the good performance of the approximate transfer function approach we will demonstrate in Section 4.
@@
    We average both sets of power spectra over a number of simulations ($100$ in our case). Note that $\tilde{C}^{{\rm PL},(\alpha\beta)(\gamma\delta)}_{b}$ does not need to be calculated empirically averaging over realisations, but instead can be alternatively calculated analytically through multiplying the input spectrum by the mask MCM, as illustrated by Equation (2.5).
@@
With the TF estimated as we just described, we can construct our model for the mode-coupling matrix of the band powers by combining it with the mask-only mode-coupling matrix (Equation 2.12). The method then proceeds as it does for the standard pseudo-$C_{\ell}$ approach: we bin and invert this mode-coupling matrix, and use it to estimate the mode-decoupled band powers and the associated BPWFs. Our implementation of this method [^6] https://github.com/simonsobs/SOOPERCOOL v0.0 includes all the cross-spectra of the $IEB$ fields. As stated, this paper will only focus on polarisation and ignore intensity.
@@ 3.1 Input simulated maps
All of our simulated maps are in healpix pixels [65], produced at single float precision. For a comparison between the numerical $-\ell$ and TF methods, shown in Section 4.1, we use maps at resolution $N_{\rm side}=128$. The limited resolution is necessary in order to compute and store the $3N_{\rm pixels}\times 3N_{\rm pixels}$ observation matrix used to incorporate the effects of filtering (see Section 3.3.2). Handling this matrix rapidly becomes challenging for resolutions $N_{\rm side}\geq 256$. For the TF method results at full resolution and multipole range, shown in Section 4.2, we use maps at resolution $N_{\rm side}=512$. All the angular power spectra in this work are calculated with NaMaster, which corrects for the mode coupling arising from partial sky observation. We do not use $E$/$B$-mode purification for simplicity. Incorporating the standard $B$-mode purification techniques that correct the $E$-to-$B$ leakage caused by the sky mask [66] is trivial to implement in this formalism (one simply replaces all fields and power spectra by their purified versions). Filtering leads to additional leakage that, ideally, should also be corrected at the map level. Methods for this exist, exploiting observation matrices to construct the covariance of the filtered data [67, 50]. This is an area of active research in SO, where the larger sky area covered makes deploying these techniques challenging. The masks are constructed from normalised hits maps and with a C1 apodisation [68] with a $5^{\circ}$ length. [^7] See https://namaster.readthedocs.io/en/latest/api/pymaster.utils.html#pymaster.utils.mask_apodization for details.
@@ 3.2 Simulated dataset
In total, our schedule has 167 observations for a total integration time of 139 hours. Each observation is almost an hour long and includes the 7 detector arrays, known as wafers, covering the focal plane of one SAT, for a total of $\sim 6000$ detectors per each f090/f150 band (however we use only one-eighth of these as described below). The normalised hits map of our full schedule is shown in Figure 2, scanning the South Galactic patch. Our schedule is run in 8 sub-schedules that produce one filter+bin map each. Sub-schedules are used to balance between the flexibility to create splits of the data and computation time. These maps are linearly co-added. [^10] Since the filtering and mapmaking operations are linear, coadding individual sub-schedule filter+bin maps yields the same map as applying the filter+bin mapmaker to the total schedule. In particular, 5 sub-schedules correspond to observations when the South patch is rising through the sky (equivalent to 102 observations and 93.5 hours) and 3 sub-schedules when the South patch is setting in the sky (equivalent to 65 observations and 45.5 hours). This separation into sub-schedules is useful when defining splits for simulating a null test or trying to estimate cross-spectra without noise bias.
@@ 3.3 Using toast
  • A HWP synchronous signal filter is used to subtract the HWP synchronous signal introduced in the modulation of the TOD. This is a sines/cosines template fitted to data as a function of HWP angle and subtracted. The HWP filter is applied to one detector and one hour-long observation at a time.
@@
  • A ground filter is used to subtract the horizontal-coordinate-fixed contamination from the ground. This is a Legendre polynomial of order 10 fitted to data as a function of azimuth and subtracted. The ground filter is applied to one detector and one hour-long observation at a time. No binning is needed: the polynomial templates are sampled at the boresight azimuth and projected out from the detector data.
  • A high-pass filter is used to suppress the impact of low-frequency contamination such as the atmosphere and detector thermal drifts, among others. This is a Legendre polynomial of order 3 fitted to the data as a function of time and subtracted. This and the ground filter are applied simultaneously since they are templates deprojected in one matrix operation. The polynomial templates are projected out separately from each sub-scan (sweep across the field).
In Appendix A, we illustrate the effect of individual filters. Finally, the TOD are binned into healpix $N_{\rm side}=512$ maps according to their pointing. We make some approximations to ease the computing cost of running these simulations for this work. We use a 40 Hz sampling rate for scanning (real observations scan at 200 Hz) and we also “thin” the focal plane by a factor of 8, i.e., only one out of every eight detectors is simulated. In Appendix B we discuss this in detail, showing how the impact is minimal while reducing the computation time by a factor of 8. The combined effect of the three data volume reductions (one day every month; 40 Hz; 1/8th of detectors) is to reduce the computational requirements of the simulations by a factor of 1200 per frequency band per SAT. With our early data analysis, we have estimated that it takes $\sim 3$M core-hours to filter the TF simulations (300 in total, 100 realisations and pure-$T$, pure-$E$ and pure-$B$ modes for each) required for one year of data with one SAT and one band. This is avoiding any data thinning, which can lead to a significant speed-up without compromising the quality of the estimated transfer function. It is therefore entirely feasible to deploy the methods described here with the high-performance computing resources available to the collaboration.